OpenStudy (anonymous):

given the coordinates of endpoint A (4, -7) and midpoint M (-2, -3) find the coordinates of B, the other endpoint in segment AB

OpenStudy (anonymous):

Find change from A to M (-6,4) add it to M we get (-8,1)

myininaya (myininaya):

\[(2(-2)-4,2(-3)-(-7))=(-4-4,-6+7)=(-8,1)\]

OpenStudy (anonymous):

\[x_{midpoint}=(1/2)(x_1+x_2)\] so: -2=(1/2)(4+x) -4=4+x -8=x \[y_{midpoint}=(1/2)(y_1+y_2)\] so -3=(1/2)(-7+y) -6=-7+y 1=y b(-8,1)

OpenStudy (anonymous):

Three different method( I think), you get to choose

myininaya (myininaya):

\[(endpointx_1,endpointy_1)=(2(midpointx)-endpointx_2,2(midpointy)-endpointy_2)\]